calculate the emission wavelengths that occur in the hydrogen spectra using the bohr equation.Hydrogen יויויין יויויויויויויוי 400 500 600 700 Fig e: Hydrogen emission spectra for calculations
Answer #1

From Bohr's theory :

Each of these lines fits following general equation, where n1 and n2 are integers and RH is 1.09678 x 10-2 nm-1.

1/ λ = RH* ( 1/n12 - 1/n22 )

As shown spectra is in visible region :   four visible hydrogen emission spectrum lines are in the Balmer series.

n1 = 2

thus,

for 2<--- 3 , 1/ λ = 1.09678 x 10-2 nm-1* ( 1/4 - 1/9 ) ==> λ = 656 nm (Red)

for 2<--- 4 , 1/ λ = 1.09678 x 10-2 nm-1* ( 1/4 - 1/16 ) ==> λ = 486 nm

for 2<--- 5 , 1/ λ = 1.09678 x 10-2 nm-1* ( 1/4 - 1/25 ) ==> λ = 434 nm

for 2<--- 6 , 1/ λ = 1.09678 x 10-2 nm-1* ( 1/4 - 1/36 ) ==> λ = 410 nm (violet)